A shopkeeper has a uniform demand of an item at the rate of 600 items per year. He buys from the supplier at a cost of Rs.8 per item. And the cost of ordering Rs. 12 each time. If the stock holding costs are 20% per year of stock value, how frequently should he replenish his stocks and what is the Optimal Order Quantity.


1
Expert's answer
2021-10-21T05:26:45-0400

We are given demand rate R=600items/year

Ordering or set up cost C3=Rs.12

Holding cost C1or C1I=Rs.8 per item

I=0.20

Economic ordering quantity (Q) ="\sqrt\frac{2C_3R}{C_1 I}"


"\sqrt\frac{2×12×600}{8×0.20}" ="\sqrt{9000}" =94.87units


Cycle time (t) ="\frac{EOQ}{R}"


"\frac{94.87}{600}=0.16years"


0.16×12=1.92months

=1.92months

Therefore the frequency that he should replenish his stocks is 1.92 months

The Optimal Order Quantity is 94.87 units